# Quantum Computing Education in Latin America: Experiences and Strategies

### v2

@@ II-C Future vision
2. Regional Collaboration: As the course proposals are initially meant to be offered online, they could reach various universities, allowing students from different institutions, cities, and even countries to access them. Expanding the scope and impact of education initiatives will involve close collaboration with universities across Latin America. We can share resources, expertise, and best practices, ensuring a cohesive educational experience for students throughout the region.
@@
3. Global Partnerships: Our vision of engaging with global quantum research communities, participating in joint projects, and facilitating student and faculty exchanges will enhance the quality of education and research, ensuring that Latin American institutions remain at the forefront of QC advancements. The courses’ funding would initially be requested from universities, but seeking support from existing organizations that promote the dissemination of QC is also an objective.
III Results and Discussion
@@
In addressing the challenge of quantum education and training, we aimed to close the gap in Latin America and the rest of the world with this first step of course creation, contributing to the unification of scientific and technological progress. We highlighted the contributions of the UNAL’s research group, an initiative that can serve as a model for other groups aiming to contribute to the field and its expansion. We also provided a practical and illustrative example of how to implement and adapt the European Competence Framework for Quantum Technologies, which not only guided the profiles and proficiency levels needed but also offered valuable information on needs, suggestions, and learning paths for each profile, useful to students once they end the course. Additionally, we introduced the new elements of “Work Teams” and “Evaluation” based on the framework to enhance the courses’ development. These additions were made possible thanks to the framework’s clarity and comprehensibility, as it offers numerous tools that allow educators to maximize its potential and create additional, valuable elements for education. Further advancement of these tools will involve examining and creating additional examples and methodologies. Once completed, these could potentially be included in a work showcasing the applications of the CFQT, as their significance lies in illustrating the learning’s applicability and unifying assessment methods, thereby contributing to a more robust and standardized approach to QC education.
@@
A critical aspect that emerged during this process was the consideration of whether to introduce QC concepts through elective courses or integrate them into mandatory foundational STEM courses. While elective courses offer flexibility and cater to specific interests, integrating quantum content into core courses ensures a broader reach across disciplines. This approach could enhance interdisciplinary learning and prepare a wider range of students for the quantum era. However, we need to take into account the short term feasibility of this approach, given that universities tend not to be very flexible regarding the modification of existing courses.
@@
Another key consideration during the course creation process was whether to offer the courses to a broad range of majors, including non-STEM students where quantum technologies are becoming increasingly relevant, or to limit them to STEM fields. We decided to make the first introductory course accessible to all students, regardless of their academic background. However, the second, more in-depth course, which focuses on more advanced quantum physics and mathematics, does have prerequisites that non-STEM students typically do not fulfill. As we move forward, it will be important to assess whether this approach is effective for engaging a diverse student population, as addressing this challenge requires innovative teaching methodologies to make quantum concepts accessible and relevant to diverse majors. Alternatively, we could explore a different strategy where STEM students interested in specific applications (e.g., cancer research) learn relevant concepts from those fields to better communicate with professionals in interdisciplinary teams. For example, QC students could gain a foundational understanding of medical concepts to facilitate collaboration with doctors, rather than expecting non-STEM professionals to master quantum concepts.
IV Conclusions
The introduction of tailored courses has the potential to significantly impact QC education across Latin America by promoting regional competitiveness and standardization. Our approach facilitates the creation of comprehensive courses that address both immediate educational needs and anticipate future advancements in QC. The framework we have adopted offers a robust foundation for standardizing QC education, enabling other universities and institutions to adapt it to their own needs. This, in turn, fosters a unified and consistent approach to quantum education and training.
@@ IV Conclusions
One of the key challenges addressed is accessibility, both in terms of geographic reach and inclusivity for students lacking enough foundations in the field. By enhancing access to QC education across Latin America, we can leverage its diverse talent pool, fostering inclusivity and a new generation of professionals who bring varied perspectives and innovative solutions to the global stage. This will not only elevate the region’s contribution to the global quantum ecosystem but also help bridge the growing skills gap worldwide, driving forward advancements in technology and its applications across various sectors. While the courses represent a step forward, further efforts are needed to ensure their scalability and adaptability. Potential limitations include relevant content selection and logistical challenges of delivering high-quality, practical QC education through online platforms.
In essence, by continuing to innovate in QC education, expanding regional and global partnerships, and focusing on reducing barriers to entry, we can ensure that the next generation of quantum scientists, engineers, and innovators is well-equipped to meet the challenges and opportunities of the quantum future.
@@ Acknowledgments
We extend our gratitude to Franziska Greinert for her feedback and insights in creating the courses. We also thank her and her team for developing the framework, which provided a comprehensive foundation for our courses design and research direction. C.E.-B. and A.M.-C. acknowledge funding from the project “Centro de Excelencia en Computación Cuántica e Inteligencia Artificial”, HERMES code $5289352893$, UNAL.

### v3

@@ II-B5 Regional Collaboration
As the course proposals are initially designed to be offered online, they have the potential to reach students from various universities, cities, and even countries across Latin America. This broader accessibility is particularly relevant in a region where quantum computing education is still developing and often concentrated in a few institutions. Expanding the scope and impact of these educational initiatives requires close collaboration between universities, research centers, and industry partners throughout Latin America.
By sharing resources, expertise, and best practices, institutions can create a more comprehensive educational experience, ensuring that students have access to high-quality quantum education, regardless of their location.
To address the language barrier, developing and translating open-access materials, as well as creating Spanish-language learning resources, could serve as a starting point to enhance inclusivity and engagement in the field. Given that Spanish is widely spoken in Latin America, these efforts would benefit a significant portion of the population, while also laying the groundwork for future initiatives that consider the region’s linguistic diversity.
Furthermore, fostering regional collaboration can help mitigate existing disparities in infrastructure and faculty training. Initiatives such as joint research projects, shared teaching materials, and inter-institutional workshops can enhance the learning experience and build a stronger quantum community within the region.
II-B6 Global Partnerships
@@
Engaging with global quantum research communities, participating in joint projects, and facilitating student and faculty exchanges will enhance the quality of education and research, providing essential opportunities for knowledge exchange, and access to cutting-edge developments and resources, while ensuring that Latin American institutions remain at the forefront of quantum advancements. The courses’ funding could initially be requested from universities, but seeking support from existing international organizations that promote the dissemination of QC is also an objective to ensure sustainability and broader impact.
II-C Outlook
As a starting point, the first proposed course is under consideration for future testing, potentially being offered online by the UNAL’s research group during a semester break. This initial phase could also involve updates to their existing Quantum Computing course available on YouTube, aligning it with clearly defined learning objectives to meet the competency requirements outlined in the framework, ensuring that the course provides a structured pathway for students to progressively develop the necessary skills. A similar approach could be valuable for other universities looking to implement the framework, as feedback from students and faculty during this phase would be crucial for refining content, delivery methods, and evaluating the effectiveness of integrating quantum education into the actual curriculum in the future, offering online and onsite courses that can be taken with the free elective credits available to students, to implement them across a broader range of careers.
Developing further undergraduate and graduate specialization courses to gain deeper understanding and practical skills in specific areas is also an objective. The courses would focus on achieving the B proficiency level and qualification profiles from P4 to P6 depending on students’ interests; as reaching the C level and the P7, P8 and P9 profiles typically requires extensive work experience, such as a PhD or post-doctoral research.
As QT continues to evolve rapidly, it is imperative to establish continuous professional development programs for educators, providing ongoing training and resources to keep instructors abreast of the latest advancements. However, in Latin America, where formal quantum education is still in its infancy, many universities lack faculty with prior expertise in QC. A potential future step in expanding QC education in the region is the development of a train-the-trainer program aimed at equipping educators with the necessary skills to teach quantum computing effectively. In addition, training courses based on the CFQT can be offered to students with more advanced knowledge, enabling them to become educators themselves. This approach not only helps increase the number of trained faculty but also facilitates the dissemination of knowledge to a broader student base, thereby expanding the reach and impact of QC education.
III Discussion
@@
In addressing the challenge of quantum education and training, we aimed to close the gap in Latin America and the rest of the world with this first step of course creation, contributing to the unification of scientific and technological progress. We highlighted the contributions of the UNAL’s research group, an initiative that can serve as a model for other groups aiming to contribute to the field and its expansion. We also provided a practical and illustrative example of how to implement the European Competence Framework for Quantum Technologies, which not only guides the profiles and proficiency levels needed but also offers valuable information on needs, suggestions, and learning paths for each profile, useful to students once they end the course.
Additionally, we introduced the elements of “Work Teams” and “Evaluation” based on the framework to enhance the courses’ development, important to illustrate applicability and unify assessment methods. These additions were made possible thanks to the framework’s clarity and comprehensibility, as it offers numerous tools that allow educators to maximize its potential and create additional elements for education. Further advancement of these tools will involve examining and creating additional examples and methodologies.
@@
A critical aspect that emerged during this process was the consideration of whether to introduce QC concepts through elective courses or integrate them into mandatory foundational STEM courses. While elective courses offer flexibility and cater to specific interests, integrating quantum content into core courses ensures a broader reach across disciplines, which could enhance interdisciplinary learning and prepare a wider range of students for the quantum era. However, we need to take into account the short term feasibility of this approach, given that universities tend not to be very flexible regarding the modification of existing courses.
@@
Another consideration during the course creation process was whether to offer the courses to a broad range of majors, including non-STEM students where quantum technologies are becoming increasingly relevant, or to limit them to STEM fields. We decided to make the first introductory course accessible to all students, regardless of their academic background. However, the second course, which focuses on more advanced quantum physics and mathematics, does have prerequisites that non-STEM students typically do not fulfill. As we move forward, it will be important to assess whether this approach is effective for engaging a diverse student population (for example by comparing the academic performance of STEM and non-STEM students, which would reveal potential gaps in understanding), as addressing this challenge requires innovative teaching methodologies to make quantum concepts accessible and relevant to diverse majors.
Alternatively, we could explore a different strategy where STEM students interested in specific applications (e.g., cancer research) learn relevant concepts from those fields to better communicate with professionals in interdisciplinary teams. For example, QC students could gain a foundational understanding of medical concepts to facilitate collaboration with doctors, rather than expecting non-STEM professionals to master quantum concepts.
IV Conclusion
This work highlights the importance of tailoring quantum education initiatives to the specific needs of Latin America while leveraging established frameworks such as the CFQT. The introduction of tailored courses has the potential to significantly impact QC education across Latin America by promoting regional competitiveness and standardization. Our approach facilitates the creation of comprehensive courses that address immediate educational needs and ensure transparent learning objectives from the outset, providing students with a clear understanding of their learning goals. Through the integration of work teams and evaluation methodologies, we demonstrate how the framework can be effectively applied to course design and assessment in the region. Additionally, we recognize that addressing the educational challenges in Latin America requires complementary strategies beyond the framework itself. By incorporating these elements, we aim to create an inclusive quantum education model that enhances accessibility and prepares students for both local and international opportunities in the field. These efforts not only contribute to strengthening quantum education in Latin America but also offer insights that may be valuable for other regions facing similar challenges.
@@ IV Conclusion
One of the key challenges addressed is accessibility, both in terms of geographic reach and inclusivity for students lacking enough foundations in the field. By enhancing access to QC education across Latin America, we can leverage its diverse talent pool, fostering a new generation of professionals who bring varied perspectives and innovative solutions to the global stage. While the courses represent a step forward, further efforts are needed to ensure their scalability and adaptability. Potential limitations include relevant content selection and logistical challenges of delivering high-quality, practical QC education through online platforms.
In essence, by continuing to innovate in quantum education, expanding regional and global partnerships, and focusing on reducing barriers to entry, we can ensure that the next generation of professionals is well-equipped to meet the challenges and opportunities of the quantum future.
@@ Acknowledgments
We extend our gratitude to Franziska Greinert for her feedback and insights in creating the courses. We also thank her and her team for developing the framework, which provided a comprehensive foundation for our courses design and research direction. C.E.-B. and A.M.-C. acknowledge funding from the project “Centro de Excelencia en Computación Cuántica e Inteligencia Artificial”, HERMES code $52893$, UNAL.